OpenStudy (joshoyen):

HELP PLS. Use the information provided to write the vertex form equation of each parabola. 1. y = -7x^2 - 14x - 6

OpenStudy (astrophysics):

Notice it's in standard form \[y=ax^2+bx+c\] right

OpenStudy (joshoyen):

yeah

OpenStudy (astrophysics):

And vertex form is \[y=a(x-h)^2+k\] where (h,k) is the vertex

OpenStudy (astrophysics):

We want it in that form

OpenStudy (astrophysics):

You can find the vertex simply by using \[\frac{ -b }{ 2a }\] so what will it be?

OpenStudy (joshoyen):

-1

OpenStudy (astrophysics):

Right, so that is our x value, now we need the corresponding y value

OpenStudy (astrophysics):

Any idea how to find it?

OpenStudy (joshoyen):

plug it in?

OpenStudy (astrophysics):

Right! Plug in for x

OpenStudy (joshoyen):

y = -7 + 14 - 6

OpenStudy (astrophysics):

Look again

OpenStudy (joshoyen):

ohhh wait

OpenStudy (joshoyen):

49 + 14 - 6

OpenStudy (astrophysics):

No you're good sorry :)

OpenStudy (joshoyen):

oh haha

OpenStudy (joshoyen):

so is that the answer?

OpenStudy (astrophysics):

So we get y = 1 right

OpenStudy (astrophysics):

Our vertex is (-1,1)

OpenStudy (astrophysics):

Now simply plug in the values into the vertex form I showed you earlier

OpenStudy (joshoyen):

wait I got y = 57

OpenStudy (astrophysics):

\(\color{blue}{\text{Originally Posted by}}\) @joshoyen y = -7 + 14 - 6 \(\color{blue}{\text{End of Quote}}\) no this is good

OpenStudy (joshoyen):

ohh okay

OpenStudy (joshoyen):

@Astrophysics wait on my paper , it says the answer is y = -7(x + 1)^2 + 1

OpenStudy (astrophysics):

Remember what I said, plug in the vertex into the vertex form I showed you earlier

OpenStudy (astrophysics):

Vertex form is \[y=a(x-h)^2+k\] where (h,k) is the vertex

OpenStudy (joshoyen):

Ohhhh, okay i see now

OpenStudy (joshoyen):

thank you

OpenStudy (astrophysics):

Yw :)